100 random samples were taken, and for each random sample we made a 95% confidence interval, about how many of those 100 confidence intervals would actually contain the parameter?

Increasing the confidence level (more than one)

a increase the width of a confidence interval

b increase the probability that the parameter is in the confidence interval

c increase the percentage of samples which will create a confidence interval that contains the parameter

d Increase the margin of error

 

A 95% confidence interval for a true mean is given by (3.3, 7.1).  What is the margin of error for this interval?

a 0.95

b 1.9

c 7.6

d 3.8
Answer #1

The number of 95% confidence intervals out of 100 would actually contain the parameter is 100 *0.95 = 95

-------------------------

Increasing the confidence level (more than one)

a increase the width of a confidence interval

c increase the percentage of samples which will create a confidence interval that contains the parameter

d Increase the margin of error

-------------------

Let \bar{x} shows the sample mean and ME shows the margin of error so

T-ME = 3.3.x + ME = 7.1

Subtracting first equation from second gives

\bar{x}+ ME-\bar{x}+ ME=7.1-3.3

2 ME=3.8

ME 1.9

Correct option is b.